# Textual Unlearning Gives a False Sense of Unlearning

### v2

@@ Abstract
1The State Key Laboratory of Blockchain and Data Security, Zhejiang University, P. R. China
2ETH Zurich, Zurich, Switzerland
@@ 1 Introduction
To address RQ1, we introduce a rigorous auditing method called Unlearning Likelihood Ratio Attack+ (U-LiRA+). By injecting and auditing the most-vulnerable samples in the training set, U-LiRA+ enables a rigorous evaluation on textual unlearning methods in the worst case, providing a precise measure of the erasing effectiveness. For RQ2, we explore the privacy risks of textual unlearning mechanisms in deployment by developing two novel attacks, collectively referred to Textual Unlearning Leakage Attacks (TULA). In the black-box scenarios, we present TULA for Membership Inference (TULA-MI), which enables an adversary to infer the membership of the unearned text just by querying the models before and after unlearning. In the white-box scenarios, we propose TULA for Data Reconstruction (TULA-DR). By transforming the text reconstruction problem into a continuous constrained-optimization task, the adversary could effectively reconstruct the unlearned texts, guided by the weights differences between original and unlearned models.
@@
(1) Unlearned texts remain detectable with high confidence on the unlearned LMs. By introducing the rigorous auditing method U-LiRA+, we argue that previous unlearning auditing methods highly overestimate the effectiveness of existing textual unlearning methods. The actual erasure effect is actually so limited that over $70\%$ of unlearned texts can still be correctly inferred, with an error rate below $0.1\%$.
(2) The textual unlearning mechanism inevitably leaks membership information of the unlearned texts. We propose TULA-MI in black-box scenarios and further analyze its variants under different levels of adversary knowledge. Our results indicate that a malicious user can confidently infer the membership of unlearned texts by comparing model outputs (even loss values) before and after unlearning.
(3) The textual unlearning mechanism poses a substantial risk of malicious reconstruction on unlearned texts. We introduce TULA-DR, enabling an adversary to reconstruct unlearned texts with even more than $80\%$ accuracy, guided by the model weights before and after unlearning.
@@ 2.1 Machine Unlearning
Machine unlearning (MU) aims to erase the impact of specific training data on a trained model (Bourtoule et al., 2021). There are generally two categories: exact unlearning and inexact unlearning. Exact unlearning methods remove the unlearned samples ($\mathcal{D}_{unlearn}$) from the training set ($\mathcal{D}$) and retrain the original model ($\mathcal{M}_{original}$) on the remaining data ($\mathcal{D}_{remian}$) from scratch, which provides the cleanest erase and is considered as the “gold standard” (Bourtoule et al., 2021). However, exact unlearning has a high computational overhead in practice, especially for LMs, given their large training corpus and massive parameters. To improve practicality and efficiency, inexact unlearning methods are proposed to efficiently approximate the retrained model without retraining from scratch. Their basic idea is to fine-tune $\mathcal{M}_{original}$ on the $\mathcal{D}_{unlearn}$ with an “inverse learning” objective. Specifically, Jang et al. propose to fine-tune $\mathcal{M}_{original}$ by Gradient Ascent (GA) based on an intuition that unlearning should be the inverse process of learning (Jang et al., 2022). Later, Chen et al. propose to maximize the Kullback-Leibler divergence of the outputs between $\mathcal{M}_{original}$ and unlearned model ($\mathcal{M}_{unlearn}$) on $\mathcal{D}_{unlearn}$ (Chen & Yang, 2023). Besides, Zhang et al. treat $\mathcal{D}_{unlearn}$ as negative preference samples and transform unlearning to a Negative Preference Optimization (NPO,  (Rafailov et al., 2024)) problem on LMs (Zhang et al., 2024).
@@ 2.2 Machine Unlearning Auditing
A successful unlearning process should ensure that the $\mathcal{M}_{unlearn}$ does not disclose any membership information indicating that $\mathcal{D}_{unlearn}$ is part of $\mathcal{D}$. Therefore, Membership Inference Attacks (MIAs) (Shokri et al., 2017) are widely used to audit MU methods (U-MIAs). MIAs aim to determine whether a sample belongs to the training set by analyzing model output discrepancies (e.g., loss values). Consequently, if the adversary cannot confidently distinguish whether a sample originates from the $\mathcal{D}_{unlearn}$ or the unseen dataset $\mathcal{D}_{unseen}$ (e.g., nearly random guess) on $\mathcal{M}_{unlearn}$, the unlearning process is deemed successful. Existing MU methods mainly use population-based U-MIAs for auditing (Shi et al., 2024), and recent works (Hayes et al., 2024; Pawelczyk et al., 2024) attempt to use per-sample U-MIAs to improve auditing precision.
@@
Per-sample U-MIAs. Per-sample U-MIA (Hayes et al., 2024; Pawelczyk et al., 2024)builds upon the Likelihood Ratio Attack (LiRA) method proposed by Carlini et al.  (Carlini et al., 2022), which formulates MIA as a hypothesis testing problem. For an audited sample $x$, the adversary models the score distributions under the hypotheses that $x$ is a member of $\mathcal{D}_{unlearn}$, and that $x$ is one of $\mathcal{D}_{unseen}$. Given the score of $x$ on $\mathcal{M}_{unlearn}$, the attack then applies a likelihood ratio test to distinguish between the two hypotheses. To estimate the distributions, the adversary trains multiple shadow models  (Shokri et al., 2017) by repeatedly sampling an identically distributed auxiliary training set $\mathcal{D}^{\prime}$. For the target $x$, the adversary trains the unseen shadow model $\mathcal{M}_{out}$ on $\mathcal{D}^{\prime}$, and trains the unlearn shadow model $\mathcal{M}_{in}$ on $x\cup\mathcal{D}^{\prime}$ after unlearning $x$. With a sufficient number of shadow models, the adversary fits two Gaussians $\mathcal{N}(\mu_{x,\mathrm{in}},\sigma_{x,\mathrm{in}}^{2})$ and $\mathcal{N}(\mu_{x,\mathrm{out}},\sigma_{x,\mathrm{out}}^{2})$ to the scores from $\mathcal{M}_{in}$ and $\mathcal{M}_{out}$ models on the target sample $x$. Finally, the adversary applies a standard Neyman–Pearson test to determine whether the observed score from the $\mathcal{M}_{unlearn}$ is more likely if $x$ is an unlearn or unseen sample: $\mathcal{A}(\mathcal{M}_{unlearn},x):=\frac{\mathcal{N}(\mathcal{S}(\mathcal{M}_{unlearn};x)\mid\mu_{x,\mathrm{in}},\sigma_{x,\mathrm{in}}^{2})}{\mathcal{N}(\mathcal{S}(\mathcal{M}_{unlearn};x)\mid\mu_{x,\mathrm{out}},\sigma_{x,\mathrm{out}}^{2})}.$
@@ 2.3 Privacy Attacks against Machine Unlearning
Chen et al. demonstrate that the MU mechanism can inadvertently reveal membership information about unlearned data (Chen et al., 2021). Specifically, for a target sample $x$, an adversary randomly selects an identically distributed auxiliary dataset $\mathcal{D}^{\prime}$ and trains two models, $\mathcal{M^{\prime}}_{original}$ and $\mathcal{M^{\prime}}_{unlearn}$ (after unlearning $x$), on $x\cup\mathcal{D}^{\prime}$. The adversary queries both models with $x$, combining their outputs to create the positive feature. For the negative feature, the adversary randomly selects another unseen sample, queries both models and combines the outputs. By repeating this process multiple times, a dataset of feature pairs is constructed to train a binary classification attack model determining whether $x$ is a member of training data. Hu et al. demonstrate that adversaries can reconstruct an unlearned sample by exploiting the difference in model weights before and after unlearning, $\Delta\theta=\mathcal{M}_{original}-\mathcal{M}_{unlearn}$ (Hu et al., 2024). This approach relies on the assumption that $\nabla^{*}$ aligns with the gradient direction of the unlearned sample $x$ on $\mathcal{M}_{original}$. Specifically, the adversary begins by inferring the label $y^{\prime}$ of the unlearned sample using a probing technique. Next, the adversary randomly initializes a dummy sample $x^{\prime}$ with the same shape as $x$ and computes the gradient $\nabla^{\prime}$ of ($x^{\prime}$,$y^{\prime}$) on $\mathcal{M}_{original}$. By minimizing the cosine similarity between $\Delta\theta$ and $\nabla^{\prime}$, the adversary iteratively updates $x^{\prime}$ until convergence. However, existing studies focus solely on image data, and the potential privacy threats against textual unlearning remain underexplored.
@@ 3 Rigorous Auditing on Textual Unlearning
Let’s go back to the drawing board to rethink the unlearning auditing. The basic idea is to test whether an adversary could confidently distinguish an unlearned sample from an unseen sample through outputs on $\mathcal{M}_{unlearn}$. However, existing auditing methods ignore a crucial and fundamental issue: the proper selection of audited samples. Typically, these methods randomly partition the dataset into unseen data $\mathcal{D}_{unseen}$ and training data $\mathcal{D}$. A subset of $\mathcal{D}$ is then randomly selected as unlearned data $\mathcal{D}_{unlearn}$, while an equal number of unseen and unlearned samples are designated as the audited samples. However, this approach inadvertently causes $\mathcal{M}_{original}$ to exhibit “false memory”, where the model acts as if it had been trained on unseen samples. Such results mainly arise from the model’s ability to generalize, as the audited samples are from the same distribution. Furthermore, the inherent overlap and redundancy among text samples exacerbate such effect (Duan et al., 2024).
The false memory effect leads to considerable overlap between the output distributions of $\mathcal{D}_{unlearn}$ and $\mathcal{D}_{unseen}$ on $\mathcal{M}_{original}$. which makes MIA-based auditing methods unable to well distinguish between the samples even before unlearning. This flawed foundation undermines the fairness and rigor of unlearning auditing, as the distinguishability between $\mathcal{D}_{unlearn}$ and $\mathcal{D}_{unseen}$ could be inherently poor.
@@ 4.1 Problem Statement
In an ML system, there are three primary roles: data contributors, model developer, and beneficiaries (including users and collaborators), as shown in Figure 1. Data contributors provide their private data to the model developer, who utilize the samples to train a model. Beneficiaries are categorized based on their access to the models. Users can query the model via the developer’s API (e.g., ChatGPT (OpenAI, 2023)). In contrast, collaborators, such as organizations contracting with the developer, have full access to the weights and can modify or deploy the model locally.
@@ 4.2.1 TULA-MI in Strict Black-Box Case
Threat Model. Given a target sample $x$, a malicious user aims to determine whether $x$ is unlearned data by querying the models before and after unlearning. In a strict black-box scenario, the user inputs $x$ via API and receives only the output score, e.g., loss value $\mathcal{L}(x,y)$, quantifying the model’s confidence on $x$. Such scenarios are common in real-world applications, such as emotion diagnosis systems that output an emotion label along with a confidence score based on dialogues. This scenario represents the minimal knowledge an adversary can access in practice, thereby indicating the lower bound of potential privacy leakage.
Method. Typically, the model yields higher confidence for training samples compared to unseen ones. For example, in terms of loss value, $\mathcal{M}_{original}$ exhibits a lower loss for an unlearned sample, while $\mathcal{M}_{unlearn}$ produces a higher one. In contrast, for an unseen sample, both models maintain similarly high loss values. This observation highlights that unlearned samples undergo significant confidences changes before and after unlearning, whereas unseen samples generally do not. Therefore, we propose TULA-MI in the strict case, leveraging confidence variations to infer the membership information of target samples. The adversary evaluates the confidence change of a target sample before and after unlearning using a threshold $\gamma$, $|\mathcal{S}(\mathcal{M}_{unlearn}(x),y)-\mathcal{S}(\mathcal{M}_{original}(x),y)|$. If the change exceeds $\gamma$, $x$ is likely to be an unlearned sample. The detailed procedure is outlined in Algorithm 4 in Appendix E.
@@ 4.2.2 TULA-MI in Relaxed Black-Box Case
Method. In this setting, a powerful adversary should be capable of performing targeted MIA for each sample. Ideally, the adversary determines the membership of $x$ by carefully fitting two distributions to its output features: the distributions when $x$ is an unlearned sample ($D^{x}_{unlearn}$) or unseen sample ($D^{x}_{unseen}$). To achieve this goal, we decompose the process into three steps: (1) Feature Space Construction: For the target sample $x$, we construct the feature vector $[l_{ori},l_{ul},l_{ori}-l_{ul}]$ based on its logits from models before and after unlearning, denoted as $l_{ori}$ and $l_{ul}$, respectively. The term $l_{ori}-l_{ul}$ serves as an augmentation to capture logit changes more effectively. (2) Distribution Estimation: We first sample a training set $\mathcal{D}$ from the auxiliary dataset, train shadow model $\mathcal{M}_{original}$ on $\mathcal{D}\cup x$, and obtain $\mathcal{M}_{unlearn}$ after unlearning $x$. The two models are then queried with $x$ to construct the output feature. By repeating this process multiple times, we estimate the $D^{x}_{unlearn}$ for $x$. For $D^{x}_{unseen}$, we sample another training set $\mathcal{D}^{\prime}$ but exclude $x$, train $\mathcal{M^{\prime}}_{original}$ on $\mathcal{D}$, and obtain $\mathcal{M^{\prime}}_{unlearn}$ after randomly unlearning a sample $x^{\prime}$. Similarly, we can obtain $D^{x}_{unseen}$ after multiple rounds of sampling and querying. (3) Attack Model Training and Inference: Features sampled from the two distributions are labeled as $1$ and 00, respectively. A classifier is then trained to infer the membership of $x$. The detailed procedure is outlined in Algorithm  2.
Algorithm 2 TULA-MI in the Relaxed Case
  Inputs: original black-box model $\mathcal{M}^{*}_{original}$, unlearned black-box model $\mathcal{M}^{*}_{unlearn}$, target sample $(x,y)$,random sample $(x,y)$, auxiliary dataset $\mathcal{D}_{aux}$, learning algorithm $A$, unlearning algorithm $U$, number of shadow models $T$, logit function $\phi$.
  $D^{x}_{unlearn}\leftarrow\{\}$, $D^{x}_{original}\leftarrow\{\}$
  while $t\leq T$ do
    $\mathcal{M}_{original}\leftarrow A(\mathcal{D}\cup(x,y))$
    $\mathcal{M}_{unlearn}\leftarrow U(\mathcal{M}_{original},(x,y)))$
    $D^{x}_{unlearn}[t]\leftarrow([l^{x}_{ori},l^{x}_{ul},l^{x}_{ori}-l^{x}_{ul}],1)$
    $\mathcal{M}^{\prime}_{original}\leftarrow A(\mathcal{D}^{\prime})$
    $D^{x}_{unseen}[t]\leftarrow([l^{x}_{ori^{\prime}},l^{x}_{ul^{\prime}},l^{x}_{ori^{\prime}}-l^{x}_{ul^{\prime}}],0)$
  end while

### v3

@@ 1 Introduction
Figure 1: Textual unlearning gives a false sense of unlearning. Our findings highlight that: (1) the unlearned data could still be detected after unlearning via rigorous auditing. (2) An adversary could instead infer the unlearned data by analyzing the models before and after unlearning.
@@
To address RQ1, we introduce a rigorous auditing method called Unlearning Likelihood Ratio Attack+ (U-LiRA+). By auditing the most-vulnerable samples in the training set (Aerni et al., 2024), U-LiRA+ enables a rigorous evaluation on the erasing effectiveness for existing unlearning methods in the worst case.
For RQ2, we explore the privacy risks of textual unlearning mechanisms in deployment by developing Textual Unlearning Leakage Attacks (TULA). In the black-box scenarios, we employ TULA for Membership Inference (TULA-MI), which enables an adversary to infer the membership of the unlearned text by querying the models before and after unlearning. In the white-box scenarios, we propose TULA for Data Reconstruction (TULA-DR). With access to model weights, the adversary could effectively reconstruct the unlearned texts via continuous constrained optimization.
@@
  • Unlearned texts remain detectable with high confidence on the unlearned LMs. We argue that previous unlearning auditing methods highly overestimate the erasing effectiveness of existing textual unlearning methods. In the worst case, over $70\%$ of unlearned texts can still be correctly inferred after unlearning, with an error rate below $0.1\%$.
  • The textual unlearning mechanism additionally introduces new risks of leaking membership information on the unlearned texts. By querying the model before and after unlearning, a malicious user can confidently infer the membership of unlearned texts by comparing the model outputs, even when only loss values are available.
  • The textual unlearning mechanism poses a substantial risk of malicious reconstruction on unlearned texts. With access to model weights before and after unlearning, an adversary could reconstruct unlearned texts with even more than $80\%$ accuracy.
Our approaches, along with the concurrent works (Pawelczyk et al., 2024; Hayes et al., 2024; Huang et al., 2024; Shumailov et al., 2024; Hu et al., 2024), highlight the importance of rigorous evaluation for unlearning mechanisms. Prior to widespread deployment, these mechanisms should undergo thorough auditing and analysis to prevent misleading conclusions and potential security risks.
@@ 2.1 Machine Unlearning
Machine unlearning aims to erase the impact of specific training data on a trained model (Bourtoule et al., 2021). There are generally two categories: exact unlearning and inexact unlearning. Exact unlearning methods remove the unlearned samples ($\mathcal{D}_{unlearn}$) from the training set ($\mathcal{D}$) and retrain the original model ($\mathcal{M}_{original}$) on the remaining data ($\mathcal{D}_{remian}$) from scratch, which provides the cleanest erase and is considered as the “gold standard” (Bourtoule et al., 2021). However, exact unlearning has a high computational overhead in practice, especially for LMs, given their large training corpus and massive parameters. To improve practicality and efficiency, inexact unlearning methods are proposed to efficiently approximate the retrained model without retraining from scratch. Their basic idea is to fine-tune $\mathcal{M}_{original}$ on the $\mathcal{D}_{unlearn}$ with an “inverse learning” objective. Specifically, Jang et al. propose to fine-tune $\mathcal{M}_{original}$ by Gradient Ascent (GA) based on an intuition that unlearning should be the inverse process of learning (Jang et al., 2022). Later, Chen et al. propose to maximize the Kullback-Leibler divergence of the outputs between $\mathcal{M}_{original}$ and unlearned model ($\mathcal{M}_{unlearn}$) on $\mathcal{D}_{unlearn}$ (Chen & Yang, 2023). Besides, Zhang et al. treat $\mathcal{D}_{unlearn}$ as negative preference samples and transform unlearning to a Negative Preference Optimization (NPO,  (Rafailov et al., 2024)) problem on LMs (Zhang et al., 2024b).
@@ 2.2 Machine Unlearning Auditing
A successful unlearning process should ensure that the $\mathcal{M}_{unlearn}$ does not disclose any membership information, indicating that $\mathcal{D}_{unlearn}$ is part of $\mathcal{D}$. Therefore, Membership Inference Attacks (MIAs) (Shokri et al., 2017) are widely used to audit MU methods (U-MIAs). MIAs aim to determine whether a sample belongs to the training set by analyzing model output discrepancies (e.g., loss values). Consequently, if the adversary cannot confidently distinguish whether a sample originates from the $\mathcal{D}_{unlearn}$ or the unseen dataset $\mathcal{D}_{unseen}$ (e.g., nearly random guess) on $\mathcal{M}_{unlearn}$, the unlearning process is deemed successful. Existing MU methods mainly use population-based U-MIAs for auditing (Shi et al., 2024), and recent works (Hayes et al., 2024; Pawelczyk et al., 2024) attempt to use per-sample U-MIAs to improve auditing precision.
@@
Per-sample U-MIAs. Per-sample U-MIA (Hayes et al., 2024; Pawelczyk et al., 2024) builds upon the Likelihood Ratio Attack (LiRA) method proposed by Carlini et al.  (Carlini et al., 2022), which formulates MIA as a hypothesis testing problem. For an audited sample $x$, the adversary models the score distributions under the hypotheses that $x$ is a member of $\mathcal{D}_{unlearn}$, and that $x$ is one of $\mathcal{D}_{unseen}$. Given the score of $x$ on $\mathcal{M}_{unlearn}$, the attack then applies a likelihood ratio test to distinguish between the two hypotheses. To estimate the distributions, the adversary trains multiple shadow models  (Shokri et al., 2017) by repeatedly sampling an identically distributed auxiliary training set $\mathcal{D}^{\prime}$. For the target $x$, the adversary trains the unseen shadow model $\mathcal{M}_{out}$ on $\mathcal{D}^{\prime}$, and trains the unlearn shadow model $\mathcal{M}_{in}$ on $x\cup\mathcal{D}^{\prime}$ after unlearning $x$. With a sufficient number of shadow models, the adversary fits two Gaussians $\mathcal{N}(\mu_{x,\mathrm{in}},\sigma_{x,\mathrm{in}}^{2})$ and $\mathcal{N}(\mu_{x,\mathrm{out}},\sigma_{x,\mathrm{out}}^{2})$ to the scores from $\mathcal{M}_{in}$ and $\mathcal{M}_{out}$ models on the target sample $x$. Finally, the adversary applies a standard Neyman–Pearson test to determine whether the observed score from the $\mathcal{M}_{unlearn}$ is more likely if $x$ is an unlearned or unseen sample:
$\mathcal{A}(\mathcal{M}_{unlearn},x):=\frac{\mathcal{N}(\mathcal{S}(\mathcal{M}_{unlearn};x)\mid\mu_{x,\mathrm{in}},\sigma_{x,\mathrm{in}}^{2})}{\mathcal{N}(\mathcal{S}(\mathcal{M}_{unlearn};x)\mid\mu_{x,\mathrm{out}},\sigma_{x,\mathrm{out}}^{2})}.$
@@ 2.3 Privacy Attacks against Machine Unlearning
Chen et al. demonstrate that the MU mechanism can inadvertently reveal membership information about unlearned data (Chen et al., 2021). Specifically, for a target sample $x$, an adversary randomly selects an identically distributed auxiliary dataset $\mathcal{D}^{\prime}$ and trains two models, $\mathcal{M^{\prime}}_{original}$ and $\mathcal{M^{\prime}}_{unlearn}$ (after unlearning $x$), on $x\cup\mathcal{D}^{\prime}$. The adversary queries both models with $x$, combining their outputs to create the positive feature. For the negative feature, the adversary randomly selects another unseen sample, queries both models, and combines the outputs. By repeating this process multiple times, a dataset of feature pairs is constructed to train a binary classification attack model determining whether $x$ is a member of training data. Hu et al. demonstrate that adversaries can reconstruct an unlearned sample by exploiting the difference in model weights before and after unlearning, $\Delta\theta=\mathcal{M}_{original}-\mathcal{M}_{unlearn}$ (Hu et al., 2024). This approach relies on the assumption that $\nabla^{*}$ aligns with the gradient direction of the unlearned sample $x$ on $\mathcal{M}_{original}$. Specifically, the adversary begins by inferring the label $y^{\prime}$ of the unlearned sample using a probing technique. Next, the adversary randomly initializes a dummy sample $x^{\prime}$ with the same shape as $x$ and computes the gradient $\nabla^{\prime}$ of ($x^{\prime}$,$y^{\prime}$) on $\mathcal{M}_{original}$. By minimizing the cosine similarity between $\Delta\theta$ and $\nabla^{\prime}$, the adversary iteratively updates $x^{\prime}$ until convergence. However, existing studies focus solely on image data, and the potential privacy threats against textual unlearning remain underexplored.
@@ 3 Rigorous Auditing on Textual Unlearning
Let’s go back to the drawing board to rethink the unlearning auditing. The basic idea is to test whether an adversary could confidently distinguish an unlearned sample from an unseen sample through outputs on $\mathcal{M}_{unlearn}$. However, existing auditing methods ignore a crucial and fundamental issue: the proper selection of audited samples. Typically, these methods randomly partition the dataset into unseen data $\mathcal{D}_{unseen}$ and training data $\mathcal{D}$. A subset of $\mathcal{D}$ is then randomly selected as unlearned data $\mathcal{D}_{unlearn}$, while an equal number of unseen and unlearned samples are designated as the audited samples. However, this approach inadvertently causes $\mathcal{M}_{original}$ to exhibit “false memory”, where the model acts as if it had been trained on unseen samples. Such results mainly arise from the model’s ability to generalize, as the audited samples are from the same distribution. Furthermore, the inherent overlap and redundancy among text samples exacerbate such an effect (Duan et al., 2024).
The false memory effect leads to considerable overlap between the output distributions of $\mathcal{D}_{unlearn}$ and $\mathcal{D}_{unseen}$ on $\mathcal{M}_{original}$, which makes MIA-based auditing methods unable to well distinguish between the samples even before unlearning. This flawed foundation undermines the fairness and rigor of unlearning auditing, as the distinguishability between $\mathcal{D}_{unlearn}$ and $\mathcal{D}_{unseen}$ could be inherently poor.
@@ 4.1 Problem Statement
In an ML system, there are three primary roles: data contributors, the model developer, and beneficiaries (including users and collaborators), as shown in Figure 2. Data contributors provide their private data to the model developer, who utilizes the samples to train a model. Beneficiaries are categorized based on their access to the models. Users can query the model via the developer’s API (e.g., ChatGPT (OpenAI, 2023)). In contrast, collaborators, such as organizations contracting with the developer, have full access to the weights and can modify or deploy the model locally.
@@ 4.2.1 TULA-MI in Strict Black-Box Case
Threat Model. Given a target sample $x$, a malicious user aims to determine whether $x$ is unlearned data by querying the models before and after unlearning. In a strict black-box scenario, the user inputs $x$ via API and receives only the output score, e.g., loss value $\mathcal{L}(x,y)$, quantifying the model’s confidence on $x$. Such scenarios are common in real-world applications, such as emotion diagnosis systems that output an emotion label along with a confidence score based on dialogues. Besides, the adversary could easily detect the occurrence of unlearning in practice by continuously querying the model with target samples. If the outputs from two consecutive sets of queries change, the adversary could execute the attack to infer whether any samples have been unlearned. The strict black-box case represents the minimal knowledge an adversary can access in practice, thereby indicating the lower bound of potential privacy leakage.
Method. Typically, the model yields higher confidence for training samples compared to unseen ones. For example, in terms of loss value, $\mathcal{M}_{original}$ exhibits a lower loss for an unlearned sample, while $\mathcal{M}_{unlearn}$ produces a higher one. In contrast, for an unseen sample, both models maintain similarly high loss values. This observation highlights that unlearned samples undergo significant confidences changes before and after unlearning, whereas unseen samples generally do not. Therefore, we propose TULA-MI in the strict case, leveraging confidence variations to infer the membership information of target samples. The adversary evaluates the confidence change of a target sample before and after unlearning using a threshold $\gamma$, $|\mathcal{S}(\mathcal{M}_{unlearn}(x),y)-\mathcal{S}(\mathcal{M}_{original}(x),y)|$. If the change exceeds $\gamma$, $x$ is likely to be an unlearned sample. The detailed procedure is outlined in Algorithm 2.
Algorithm 2 TULA-MI in the Strict Case
  Inputs: Black-box models $\mathcal{M}_{original}$, $\mathcal{M}_{unlearn}$, target sample $(x,y)$, score function $\mathcal{S}$,threshold $\gamma$.
  if  $|\mathcal{S}(\mathcal{M}_{unlearn}(x),y)-\mathcal{S}(\mathcal{M}_{original}(x),y)|>\gamma$  then
    Return $(x,y)$ is an unlearned sample
  else
    Return $(x,y)$ is an unseen sample
  end if
@@ 4.2.2 TULA-MI in Relaxed Black-Box Case
Method. In this setting, a powerful adversary should be capable of performing targeted MIA for each sample. Ideally, the adversary determines the membership of $x$ by carefully fitting two distributions to its output features: the distributions when $x$ is an unlearned sample ($D^{x}_{unlearn}$) or an unseen sample ($D^{x}_{unseen}$). To achieve this goal, we decompose the process into three steps: (1) Feature Space Construction: For the target sample $x$, we construct the feature vector $[l_{ori},l_{ul},l_{ori}-l_{ul}]$ based on its logits from models before and after unlearning, denoted as $l_{ori}$ and $l_{ul}$, respectively. The term $l_{ori}-l_{ul}$ serves as an augmentation to capture logit changes more effectively. (2) Distribution Estimation: We first sample a training set $\mathcal{D}$ from the auxiliary dataset, train shadow model $\mathcal{M}_{original}$ on $\mathcal{D}\cup x$, and obtain $\mathcal{M}_{unlearn}$ after unlearning $x$. The two models are then queried with $x$ to construct the output feature. By repeating this process multiple times, we estimate the $D^{x}_{unlearn}$ for $x$. For $D^{x}_{unseen}$, we sample another training set $\mathcal{D}^{\prime}$ but exclude $x$, train $\mathcal{M^{\prime}}_{original}$ on $\mathcal{D}$, and obtain $\mathcal{M^{\prime}}_{unlearn}$ after randomly unlearning a sample $x^{\prime}$. Similarly, we can obtain $D^{x}_{unseen}$ after multiple rounds of sampling and querying. (3) Attack Model Training: Features sampled from the two distributions are labeled as $1$ and 00, respectively. A classifier is then trained to infer the membership of $x$. The detailed procedure is outlined in Algorithm  4 in Appendix G.